# Maintaining Adversarial Robustness in Continuous Learning

### v1

@@ A.2 Matrix Composition
A.3 Reshape Prior to Performing SVD
@@
Figure 7: (a) Performing SVD on an example $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ with the shape $(n,c_{l}h_{l}w_{l})$ obtains the base vectors that can constrain the updates $\Delta\widetilde{\mathbf{W}}^{l}$. Here, $x_{i,j}^{l}$ denotes the $i$-th feature of $j$-th input of $l$-th layer, and an single input $\mathbf{x}^{l}$ from $\mathbf{X}^{l}$ is illustrated on the left of Fig. 6. (b) The orthogonality between any base vector and each $h_{l+1}w_{l+1}(=4)$ column vectors of $\widetilde{\mathbf{W}}^{l}$ (see the right of Fig.6) is equivalent to the orthogonality between $h_{l+1}w_{l+1}$ sub-vectors of the base vector and the weight $\mathbf{w}^{l}$ of the kernel. (c) Prior to performing SVD, each row vector in $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ is reshaped into a matrix consisting of $h_{l+1}w_{l+1}$ row vectors with a length of $c_{l}k_{l}k_{l}$. Consequently, the shape of $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ is modified to $(nh_{l+1}w_{l+1},c_{l}k_{l}k_{l})$.
@@ B.3 Hyper-parameter Configurations
Adversarial attack method method.
@@
Table 4: Hyper-parameter setup for different adversarial attacks
Continuous learning method. We run the methods of SI, EWC, GEM, A-GEM based on the Avalanche [9], an end-to-end continual learning library. In DGP, $\alpha_{1}$ and $\alpha_{2}$ (see $\alpha$ in Eq. 7 of main text) control the number of base vectors added into the pool for stabilizing the final output and sample gradients respectively. $\alpha_{3}$ is used in reducing the number of base vectors when the pool is full (by performing the SVD and $k$-rank approximation on the matrix consisting of all base vectors in the pool).
@@
GPM memory_size $=100$, $*\alpha 1=0.97+0.003*$task_id
@@
memory_size $=100$, $*\alpha 1=0.97+0.003*$task_id,
@@
GPM memory_size $=100$, $*\alpha 1=0.985+0.003*$task_id
@@
Table 5: Hyper-parameter setup in our approach and baselines.

### v2

@@ A.2 Matrix Composition
A.3 Reshape $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ Prior to Performing SVD
@@
Figure 7: (a) Performing SVD on an example $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ with the shape $(n,c_{l}h_{l}w_{l})$ obtains the base vectors that can constrain the updates $\Delta\widetilde{\mathbf{W}}^{l}$. Here, $x_{i,j}^{l}$ denotes the $i$-th feature of $j$-th input of $l$-th layer, and an single input $\mathbf{x}^{l}$ from $\mathbf{X}^{l}$ is illustrated on the left of Fig. 6. (b) The orthogonality between any base vector and each $h_{l+1}w_{l+1}(=4)$ column vectors of $\widetilde{\mathbf{W}}^{l}$ (see the right of Fig.6) is equivalent to the orthogonality between $h_{l+1}w_{l+1}$ sub-vectors of the base vector and the weight $\mathbf{w}^{l}$ of the kernel. (c) Prior to performing SVD, each row vector in $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ is reshaped into a matrix consisting of $h_{l+1}w_{l+1}$ row vectors with a length of $c_{l}k_{l}k_{l}$. Consequently, the shape of $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ is modified to $(nh_{l+1}w_{l+1},c_{l}k_{l}k_{l})$. This results that the base vectors obtained from performing SVD on $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ have the same shape $\left(c_{l}k_{l}k_{l},\right)$ with the flattened convolution kernel in the $l$-th layer, and can be directly used to constrain the weight updates of the convolution kernels.
@@ A.4 Gradients in Batch Normalization Layer
A.5 Computational Complexity Increment
Compared to the naive training procedure, i.e., SGD, the increase of training complexity in the proposed method is mainly related to SVD. After finishing the training on each new task, we gather the layer-wise outputs and their gradient with respect to samples, and perform SVD on them to obtain the base vectors used for gradient projection. Specifically, we call the interface in Pytorch to perform SVD decomposition. This interface uses the Jacobi method with a time complexity of approximately ${o}\left(nm_{l}\min\left(n,m_{l}\right)\right)$, where $n$ is the sample number and $m_{l}$ is the features number of the $l$-th layer’s output. Assuming that the neural network consists of $L$ layers, with each layer’s outputs having an equal number of features $m$, the proposed method introduces a computational complexity increment of ${o}\left(Lnm_{l}\min\left(n,m_{l}\right)\right)$.
@@ B.3 Hyper-parameter Configurations
B.3.1 Adversarial attack algorithm
The norm of attacks used in experiments are $\ell_{\infty}$. The hyper-parameters in various attack algorithm are provided in Table. 4.
@@ B.3.1 Adversarial attack algorithm
Table 4: Hyper-parameter setup to control the attack strength
The perturbation size used in our experiments are smaller than the typical value in adversarial robustness literature. This adjustment is made because when confronted with such intensity of adversarial attacks, regardless of approaches considered (including baselines and the proposed method), the neural network’s robustness on the current task decreases significantly after learning only two or three news tasks. Thus, we slightly reduced the perturbation size. Then, the advantage of the proposed method becomes evident. While most baselines still exhibit a significant decrease after learning only two or three new tasks, the proposed method enables maintaining the model’s robustness after learning a sequence of new tasks.
B.3.2 Continual learning algorithm
We outline the fundamental principle for each continual learning algorithm in baselines as follows:
  • $\bullet$
    EWC is a regularization technique that utilizes the Fisher Information matrix to quantify the contribution of model parameters on preserving knowledge of previous tasks;
  • $\bullet$
    SI computes the local impact of model parameters on global loss variations, consolidating crucial synapses by preventing their modification in new tasks;
  • $\bullet$
    A-GEM is a memory-based approach, similar to GEM, which leverages data from episodic memory to adjust the gradient direction of current model update;
  • $\bullet$
    OGD is another gradient projection approach where each base vector constrains the weight updates of the entire model, while GPM employs a layer-wise gradient projection strategy.
We run the methods of SI, EWC, GEM, A-GEM based on the Avalanche [11], an end-to-end continual learning library. In DGP, $\alpha_{1}$ and $\alpha_{2}$ (see $\alpha$ in Eq. 6 of main text) control the number of base vectors added into the pool for stabilizing the final output and sample gradients respectively. $\alpha_{3}$ is used in reducing the number of base vectors when the pool is full (by performing the SVD and $k$-rank approximation on the matrix consisting of all base vectors in the pool).
@@ B.3.2 Continual learning algorithm
GPM memory_size $=100$, $\alpha 1=0.97+0.003*$task_id
@@
memory_size $=100$, $\alpha 1=0.97+0.003*$task_id,
@@
GPM memory_size $=100$, $\alpha 1=0.985+0.003*$task_id
@@
Table 5: Hyper-parameter setup in our approach and other CL algorithms in baselines.
B.4 Incompatibility Between Existing Continual Learning and Defense Algorithms
Distillation is a well-known defense method [25] that involves training two models - a teacher model is trained using one-hot ground truth labels and a student model is trained using the softmax probability outputs of the teacher model. The result of combinations of Distillation and existing continual learning algorithms are presented in Fig. 8. There is a notable trend in Fig.8d: the blue line (representing the performance of Distill+GPM on original samples) exhibits a more rapid decline compared to the corresponding blue line in the fourth subplot of the first row in Fig.3 (representing IGR+GPM), as well as the pink line in Fig.4d (representing AT+GPM). Additionally, the purple and blue lines in Fig.8d (representing Distill+GEM and Distill+GPM) closely align with the green line in Fig.8d (representing Distill+SGD). These observations suggest again incorporating the defense algorithms, such as Distillation, into the training procedure compromise the efficacy of these continual learning methods.
Figure 8: As Fig. 3, but for defense algorithms Distillation on PMNIST dataset. Here, we combine Distillation with continual learning methods GEM and GPM, which have shown superior ACC compared to other baselines in Fig. 3.